OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through (2, -1) and parallel to a line with slope of 3/4 help @RolyPoly

OpenStudy (anonymous):

Parallel lines have the same slope so the slope, m, of this line would be 3/4 y-y1=m(x-x1) y-(-1)=3/4(x-2) y+1=3/4(x-2) Multiply both sides of the equation by 4 4y+4=3(x-2) 4y+4=3x-6 3x-4y=-10

OpenStudy (anonymous):

since the line is paralell, the slope is the same the equation of the new line will use the point (2,-1) and m=3/4 (y-(-1))=(3/4)(x-2) simplify this to y+1=(3/4)x-(3/2) in standard form, get the x and y to one side, and the constants to the other y-(3/4)x=(-3/2)-1 y-(3/4)x=(-5/2) eliminate fractions by multiplying everything by 4 4y-3x=-10

OpenStudy (anonymous):

General form is ax+by=c Where a, b and c are integers whole numbers, natural numbers and their additive inverses like...(-1,0,1)

OpenStudy (anonymous):

thanks!!:D

OpenStudy (anonymous):

Does your book or teacher represent the general form of an equation like this? Because general form and standard form are often intertwined it seems... If so, take note of the difference between general form and standard form [To my knowledge standard form of an equation of a line would be ax+by-c=0]

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@purplec16 it appears that you math was correct up until 4y+4=3x-6 3x-4y=-10 the 4y was subtracted properly, but the 6 was not properly added to the other side the answer should have been 3x-4y=+10 not 3x-4y=-10

OpenStudy (anonymous):

Oh my apologies @justpicksomethin I should have not crossed over in my head at 1:00 a.m.